Find an equation of the line having the given slope and containing the given point. m=6,(7,1)
the equation of the line in slope-intercept is y=

Answer by ankor@dixie-net.com(22740)

Find an equation of the line having the given slope and containing the given point. m=6,(7,1)
:
Start with the point/slope form: y - y1 = m(x - x1)
y - 1 = 6(x - 7)
y - 1 = 6x - 42
y = 6x - 42 + 1
:
the equation of the line in slope-intercept is y = 6x - 41